OpenStudy (anonymous):

(1/x)-(1/y)/(y/x^2)-(1/y)

OpenStudy (campbell_st):

rewrite the numerator with a common denominator 1/x - 1/y = (y - x)/(xy) similar process for the denominator the the rule for dividing by a fraction is flip and multiply look for common factors that can be cancelled.

OpenStudy (anonymous):

So I rewrite the numerator and the denominator?

OpenStudy (campbell_st):

yes.... the denominator is y/x^2 - 1/y = (y^2-x^2)/(x^2y) the numerator in this part is the difference of 2 squares... and can be factored. once you have done that, flip it then multiply by the numerator shown above you should see some common factors that cancel